# Who Pays for Fairness? Rethinking Recourse under Social Burden

### v1

@@ 5.2 Results
Moreover, the results show that improving fairness at the prediction level does not guarantee fairer recourse outcomes and can even exacerbate them. In particular, enforcing parity in TPR through POSTPRO significantly increases the burden and cost for decision subjects. This occurs when prediction-level fairness improvements are misaligned with fairness improvements at recourse. On the contrary, the holistic perspective of MISOB enables effective improvements in group TPR, while reducing cost and burden of individuals, underscoring the importance of considering the entire pipeline. Besides, note that POSTPRO requires training a separate model for each sensitive group characterization, whereas MISOB, being blind to sensitive information, needs to be trained only once and can be evaluated across multiple group characterizations.
The effect of $C$.
@@
The performance and fairness of MISOB are governed by the emphasis placed on high-burden individuals, controlled by the weighting parameter $C$ from Equation 4, as it can be seen in Figure 2. Increasing $C$ (i.e., the importance of high-burden instances) strengthens minimax fairness across both prediction and recourse by improving worst-group metrics: TPR and AR increase, while cost and burden decrease. However, excessively large values of $C$ can reduce overall accuracy due to the overemphasis on high-burden cases. Nonetheless, there is a broad range of $C$ values (e.g., $[0.1,0.5]$ in Figure 2(a)) where MISOB achieves substantial fairness gains with minimal impact on accuracy. Thus, $C$ governs the trade-off between fairness and accuracy, allowing users to select the value that best aligns with their objectives. In our experiments, we use $C=0.3$, which achieves strong fairness gains without compromising overall performance.
Figure 2: The performance and fairness guarantees of MISOB for different values of $C$ to compute the weighted loss in the training process for recourse methods (left) GS and (right) WT, evaluated on the Adult dataset with (a,b) race and (c,d) gender as the sensitive attributes. Increasing $C$, which enhances the emphasis on high-burden instances, strengthens minimax fairness guarantees. However, very large $C$ values may hurt overall accuracy.
@@ 6 Discussion and Limitations
We now discuss some limitations of our approach that open up further avenues for future work. Our work introduces a novel framework to compute fairness metrics in a recourse-aware setting and empirically validates the effectiveness of MISOB. Analyzing the convergence properties of the training procedure (e.g., identifying stationary points) would be a valuable direction for further research. One approach might be to frame recourse fairness as a two-player Stackelberg game, following the strategic classification perspective [8]. Additionally, applying calibration techniques [21] could improve prediction reliability (cf. Equation 3), which in turn may further enhance fairness when combined with MISOB. Finally, fairness concerns may emerge over time due to model updates, dataset shifts [7] or prediction performativity [32, 39], and recourse recommendations have been shown to be temporally fragile [12]. Extending our approach to measure and optimize the social burden of recourse in dynamic settings is, therefore, a natural path for future research. Furthermore, we aim to explore more realistic cost models that better reflect the complexities of the real world and assess how these influence fairness guarantees [46].
@@ 7 Conclusions
In this work, we study the unfairness that can arise in automated decision-making pipelines that provide recourse. We provide a theoretical characterization of the sources of unfairness in algorithmic recourse, formally linking fairness guarantees in recourse and prediction, and further highlight the limitations of the standard equal cost paradigm. To address these challenges, we introduce MISOB, a novel framework grounded in the notion of social burden, which enables a holistic treatment of fairness and goes beyond a gap-based characterization of fairness guarantees. MISOB is broadly applicable: it is compatible with any classifier–recourse pair and does not require access to sensitive attributes during training or inference. Empirical results show that improving fairness solely at the prediction level is often insufficient, and can even considerably increase the cost and burden of decision subjects. In contrast, our holistic approach leads to consistent joint fairness improvements in prediction and recourse, without compromising overall performance.
Acknowledgements
The work was partially supported by the following projects: Horizon Europe Programme, grants #101120237-ELIAS and #101120763-TANGO. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Proofs of Section 3
@@ A.1 Proof of Equation 2
Proof.
The expected recourse cost for group $\boldsymbol{s}$, computed over the entire decision pipeline (i.e., accounting for how the full population is treated throughout the process) is formally defined as:
$\mathcal{C}_{f,g}^{\boldsymbol{s}}=\mathbb{E}_{p(X|S=\boldsymbol{s})}\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}$ (A.1)
The above expression can be decomposed into the contributions from instances receiving positive predictions and those receiving negative predictions as follows:
$\mathcal{C}_{f,g}^{\boldsymbol{s}}=\mathbb{E}_{p(X|S=\boldsymbol{s},f(\boldsymbol{x})=0)}\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}\cdot P(f(\boldsymbol{x})=0|S=\boldsymbol{s})$ (A.2)
$+\mathbb{E}_{p(X|S=\boldsymbol{s},f(\boldsymbol{x})=1)}\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}\cdot P(f(\boldsymbol{x})=1|S=\boldsymbol{s})$
We assume that those instances that have received a positive classification (i.e., $\boldsymbol{x}:f(\boldsymbol{x})=1$) do not change their features and, therefore, $\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))=0$ for those instances. Consequently, the second term vanishes and since $P(f(\boldsymbol{x})=0|S=\boldsymbol{s})+P(f(\boldsymbol{x})=1|S=\boldsymbol{s})=1$, we get
$\mathcal{C}_{f,g}^{\boldsymbol{s}}=\mathbb{E}_{p(X|S=\boldsymbol{s},f(\boldsymbol{x})=0)}\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}\cdot(1-P(f(\boldsymbol{x})=1|S=\boldsymbol{s}))$ (A.3)
@@ A.2 Proof of Equation 3
Proof.
The concept of social burden analyses the effect of the algorithmic pipeline on the instances that have a true positive label:
$\mathcal{B}_{f,g}^{\boldsymbol{s}}=\mathbb{E}_{p(X|S=\boldsymbol{s},Y=1)}\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}$ (A.4)
The former expression can be decomposed as follows:
$\mathcal{B}_{f,g}^{\boldsymbol{s}}=\mathbb{E}_{p(X|S=\boldsymbol{s},Y=1,f(\boldsymbol{x})=0)}$ $\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}$ (A.5)
$\cdot P(f(\boldsymbol{x})=0|S=\boldsymbol{s},Y=1)$
$+\mathbb{E}_{p(X|S=\boldsymbol{s},Y=1,f(\boldsymbol{x})=1)}$ $\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}$
$\cdot P(f(\boldsymbol{x})=1|S=\boldsymbol{s},Y=1)$
We assume that those individuals that have received a positive classification (i.e., $\boldsymbol{x}:f(\boldsymbol{x})=1$) do not change their features and, therefore, $\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))=0$ for those instances. Consequently, the second term vanishes and since $P(f(x)=0|S=\boldsymbol{s},Y=1)+P(f(x)=1|S=\boldsymbol{s},Y=1)=1$, we get
$\mathcal{B}_{f,g}^{\boldsymbol{s}}=$ $\mathbb{E}_{p(X|S=\boldsymbol{s},Y=1,f(\boldsymbol{x})=0)}\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}\cdot$
$(1-P(f(\boldsymbol{x})=1|S=\boldsymbol{s},Y=1))$
Appendix B Detailed Experimental Setting
@@ B.1 Datasets
It is a dataset based on the data from the 1994 US Census, where the main goal is to predict whether an individual earns more than 50,000\$ per year. The 14 features used to describe the instances include occupation, marital status, and education. Furthermore, it contains sensitive information such as age, gender, and race. In our experiments, we consider sensitive groups based on gender and race, as well as intersectional groups defined by the combination of these two attributes. This dataset is publicly available in the UCI repository. [^6] http://archive.ics.uci.edu/ml/index.php Although the dataset is originally divided into predefined training and test sets, we merge them and generate new train/test splits for our empirical evaluation. We pre-process the dataset as detailed in Pawelczyk et al. [38].
@@
This dataset [^7] https://www.kaggle.com/c/GiveMeSomeCredit/ includes financial and demographic information for 150,000 individuals and is used to predict the likelihood of experiencing financial distress within two years. Each record consists of 10 feature variables, such as age, monthly income, and credit indicators, and a binary target variable indicating whether the individual experienced a serious delinquency (90+ days past due) within that timeframe (0 = Yes, 1 = No). For our fairness-aware analysis, we used age as the sensitive attribute, dividing individuals into two groups: those aged 30 or younger, and those older than 30. This distinction reflects real socioeconomic and financial differences relevant to credit risk modeling. Younger individuals often lack established credit histories and income stability, whereas older individuals tend to have more financial experience and responsibilities, which can affect their credit behavior in meaningful ways. Moreover, in our experiments, we applied the following modifications:
@@
    We pre-processed the dataset following the procedure of De Toni et al. [10] to ensure recourse tractability.
@@
The German Credit dataset collects information about several individuals created from a German bank’s data from 1994. It contains details about the socioeconomic situation of individuals: their employment, housing, savings, etc. Besides, the set of features includes some sensitive information as well, such as gender and age. In this classification task, the objective is to predict whether an individual should obtain a good or bad credit score. This dataset is considerably smaller than the previous ones, containing only 1,000 instances and 20 features, and it is publicly available in the UCI repository. In our experiments, age has been considered as the sensitive attribute, and we have divided the instances into two (sensitive) groups based on whether their age is greater than 30 or less than or equal to 30. We pre-processed the dataset as detailed in [38].

### v3

@@ Results.
Table 2 shows that training the base classifier under the MISOB framework improves both predictive and recourse-related performance metrics across all groups, without compromising overall accuracy. Notably, MISOB enhances worst-case group outcomes and, in many cases, reduces disparities across groups, achieving fairness improvements without degrading the performance for privileged groups. These results also highlight an important limitation of gap-based fairness metrics: small gaps can mask universally poor outcomes across groups, giving a misleading sense of fairness. In contrast, MISOB can produce configurations with slightly larger gaps but better metric values for every group, thereby promoting greater overall well-being. Moreover, the results show that improving fairness at the prediction level does not guarantee fairer recourse outcomes and can even exacerbate them. In particular, enforcing parity in TPR through POSTPRO significantly increases the burden and cost for decision subjects. This occurs when prediction-level fairness improvements are misaligned with fairness improvements at recourse. On the contrary, the holistic perspective of MISOB enables effective improvements in group TPR, while reducing the cost and burden of individuals, underscoring the importance of considering the entire pipeline. Besides, note that POSTPRO requires training a separate model for each sensitive group characterization, whereas MISOB, being blind to sensitive information, needs to be trained only once and can be evaluated across multiple group characterizations.
@@
The performance and fairness of MISOB are governed by the emphasis placed on high-burden individuals, controlled by the weighting parameter $\alpha$ from Eqn. 4. Fig. 2 shows that increasing $\alpha$ (i.e., the importance of high-burden instances) strengthens minimax fairness across both prediction and recourse by improving worst-group metrics. However, larger values of $\alpha$ ($\alpha>0.5$) can reduce overall accuracy due to the overemphasis on high-burden cases. In practice, $\alpha$ governs the trade-off between fairness and accuracy, allowing users to select the value that best aligns with their objectives.
@@ 6 Discussion and Limitations
Our work introduces a novel framework to compute fairness metrics in a recourse-aware setting and empirically validates the effectiveness of MISOB. Analyzing the convergence properties of the training procedure (e.g., identifying stationary points) would be a valuable direction for further research. One approach might be to frame recourse fairness as a two-player Stackelberg game, following the strategic classification perspective (chen2023learning). Additionally, applying calibration techniques (guo2017calibration) could improve prediction reliability (cf. Eqn. 3), which in turn may further enhance fairness when combined with MISOB. Finally, further fairness concerns may emerge due to model updates, dataset shifts (castelnovo2021towards), prediction performativity (liu2018delayed; perdomo2020performative), or recourse recommendations that are not robust over time (detoni2025temporal). Extending our approach to measure and optimize the social burden of recourse in dynamic settings is, therefore, a natural path for future research. Lastly, we aim to explore more realistic cost models that better reflect the real world’s complexities and assess how these influence fairness guarantees (tominaga2024reassessing; esfahani2024preference).
@@ 7 Conclusions
In this work, we study the unfairness that can arise in automated decision-making pipelines that provide recourse. We provide a theoretical characterization of the sources of unfairness in algorithmic recourse, formally linking fairness guarantees in recourse and prediction, and further highlight the limitations of the standard equal cost paradigm. To address these challenges, we introduce MISOB, a novel framework grounded in the notion of social burden, which enables a holistic treatment of fairness and goes beyond a gap-based characterization of fairness guarantees. MISOB is broadly applicable: it is compatible with many classifier–recourse pairs and does not require access to sensitive attributes during training or inference. Empirical results show that improving fairness solely at the prediction level is often insufficient and can even considerably increase the cost and burden of decision subjects. In contrast, our approach leads to consistent joint fairness improvements in prediction and recourse, without compromising overall performance.
Acknowledgments
We would like to thank all the reviewers for their insightful feedback. We also thank Mikel Malagon for the technical support and for reading preliminary versions of the paper. This research was funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. This work is supported by the European Research Council under the European Union’s Horizon 2020 research and innovation programme Grant Agreement no. 851538 - BayesianGDPR, Horizon Europe research and innovation programme Grant Agreement no. 101120763 - TANGO, and no. 101120237 - ELIAS. This work is also supported by the Basque Government under grant IT1504-22 and through the BERC 2022-2025 program; by the Spanish Ministry of Science and Innovation under the grant PID2022-137442NB-I00, and through BCAM Severo Ochoa accreditation CEX2021-001142-S / MICIN / AEI / 10.13039/501100011033. Lastly, this work was also supported by Ministero delle Imprese e del Made in Italy (IPCEI Cloud DM 27 giugno 2022 – IPCEI-CL-0000007), funded by the European Commission under the NextGeneration EU programme.
Appendix
Appendix A Proofs
@@ A.1 Proof of Equation 2
Let us denote the recourse cost with the random variable $Z=\delta((X,S),g_{f}(X))$. Further, we denote with $P(f(X)=0\mid S=s)$ the likelihood of being classified negatively, conditioned on the protected group $s\in\mathcal{S}$. Then, the expected recourse cost for a given protected group $s\in\mathcal{S}$ can be defined as follows:
$\mathbb{E}[Z\mid S=s]$ $=\sum_{x\in\mathcal{X}}Z(x,s)P(X=x\mid S=s)$ (10)
$=\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)P(X=x^{\prime}\mid S=s)+\sum_{x^{\prime\prime}\in\mathcal{X}:f(x^{\prime\prime})=1}Z(x^{\prime\prime},s)P(X=x^{\prime\prime}\mid S=s)$ (11)
$\overset{(i)}{=}\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)P(X=x^{\prime}\mid S=s)$ (12)
$=\frac{P(f(X)=0\mid S=s)}{P(f(X)=0\mid S=s)}\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)P(X=x^{\prime}\mid S=s)$ (13)
$=P(f(X)=0\mid S=s)\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)\frac{P(X=x^{\prime}\mid S=s)}{P(f(X)=0\mid S=s)}$ (14)
$=P(f(X)=0\mid S=s)\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)\frac{P(X=x^{\prime}\mid S=s)}{\sum_{\mathcal{X}}\mathds{1}\{f(x^{\prime\prime\prime})=0\}P(X=x^{\prime\prime\prime}\mid S=s)}$ (15)
$=P(f(X)=0\mid S=s)\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)\frac{P(X=x^{\prime}\mid S=s)}{\sum_{x^{\prime\prime\prime}\in\mathcal{X}:f(x^{\prime\prime\prime})=0}P(X=x^{\prime\prime\prime}\mid S=s)}$ (16)
$=P(f(X)=0\mid S=s)\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)P(X=x^{\prime}\mid S=s,f(X=x^{\prime})=0)$ (17)
$=P(f(X)=0\mid S=s)\mathbb{E}[Z\mid S=s,f(X)=0]$ (18)
$=\mathbb{E}[Z\mid S=s,f(X)=0](1-P(f(X)=1\mid S=s)$ (19)
where $(i)$ arises from assuming that those instances that have received a positive classification (i.e., $x^{\prime}\in\mathcal{X}:f(x^{\prime})=1$) do not change their features and, therefore, $\delta((x,s),g_{f}(x))=0$. Then, by substituting $Z$, we arrive at the expected recourse cost:
$C_{f,g}^{s}$ $=\mathbb{E}[Z\mid S=s,f(X)=0](1-P(f(X)=1\mid S=s)$ (20)
$=\mathbb{E}[\delta((X,S=s),g_{f}(X))\mid S=s,f(X)=0](1-P(f(X)=1\mid S=s))$ (21)
In summary, the actual expected cost of algorithmic recourse depends on two factors: (i) the effort required of individuals in the sensitive group $s\in\mathcal{S}$ who receive a negative classification, quantified by $\mathbb{E}[\delta((X,S=s),g_{f}(X))\mid S=s,f(X)=0]$, and (ii) the probability that these individuals receive a negative outcome, $1-P(f(X)=1\mid S=s)$.
@@ A.2 Proof of Equation 3
We follow a similar setup to the one in Section A.1. Let us denote the recourse cost with the random variable $Z=\delta((X,S),g_{f}(X))$. We denote with $P(f(X)=0\mid S=s,Y=1)$ the likelihood of being incorrectly classified negatively by the model, conditioned on the protected group $s\in\mathcal{S}$, given that the ground truth outcome is positive $Y=1$. Then, the expected social burden for a given protected group $s\in\mathcal{S}$ can be defined similarly as in Section A.1. We report the full proof here for completeness:
$\mathbb{E}[Z\mid S=s,Y=1]$ $=\sum_{x\in\mathcal{X}}Z(x,s)P(X=x\mid S=s,,Y=1)$ (22)
$=\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)P(X=x^{\prime}\mid S=s,Y=1)+\sum_{x^{\prime\prime}\in\mathcal{X}:f(x^{\prime\prime})=1}Z(x^{\prime\prime},s)P(X=x^{\prime\prime}\mid S=s,Y=1)$ (23)
$\overset{(i)}{=}\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)P(X=x^{\prime}\mid S=s,Y=1)$ (24)
$=\frac{P(f(X)=0\mid S=s,Y=1)}{P(f(X)=0\mid S=s,Y=1)}\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)P(X=x^{\prime}\mid S=s,Y=1)$ (25)
$=P(f(X)=0\mid S=s,Y=1)\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)\frac{P(X=x^{\prime}\mid S=s,Y=1)}{P(f(X)=0\mid S=s,Y=1)}$ (26)
$=P(f(X)=0\mid S=s,Y=1)\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)\frac{P(X=x^{\prime}\mid S=s,Y=1)}{\sum_{\mathcal{X}}\mathds{1}\{f(x^{\prime\prime\prime})=0\}P(X=x^{\prime\prime\prime}\mid S=s,Y=1)}$ (27)
$=P(f(X)=0\mid S=s,Y=1)\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)\frac{P(X=x^{\prime}\mid S=s,Y=1)}{\sum_{x^{\prime\prime\prime}\in\mathcal{X}:f(x^{\prime\prime\prime})=0}P(X=x^{\prime\prime\prime}\mid S=s,Y=1)}$ (28)
$=P(f(X)=0\mid S=s,Y=1)\sum_{x^{\prime}\in\mathcal{X}:f(x^{\prime})=0}Z(x^{\prime},s)P(X=x^{\prime}\mid S=s,f(X=x^{\prime})=0,Y=1)$ (29)
$=P(f(X)=0\mid S=s,Y=1)\mathbb{E}[Z\mid S=s,f(X)=0,Y=1]$ (30)
$=\mathbb{E}[Z\mid S=s,f(X)=0,Y=1](1-P(f(X)=1\mid S=s,Y=1))$ (31)
where $(i)$ arises from assuming that those instances that have received a positive classification (i.e., $x^{\prime}\in\mathcal{X}:f(x^{\prime})=1$) do not change their features and, therefore, $\delta((x,s),g_{f}(x))=0$. Then, by substituting $Z$, we arrive at the expected social burden:
$B_{f,g}^{s}$ $=\mathbb{E}[Z\mid S=s,f(X)=0,Y=1](1-P(f(X)=1\mid S=s,Y=1))$ (32)
$=\mathbb{E}[\delta((X,S=s),g_{f}(X))\mid S=s,f(X)=0,Y=1](1-P(f(X)=1\mid S=s,Y=1))$ (33)
Appendix B Further Implementation Details for the Experiments
@@ B.1 Datasets
In the paper, we used three fairly common datasets both in the recourse and fairness literature. Specifically, we used Adult (adult_2) for the experiments in Section 5, and GiveMeSomeCredit [^7] https://www.kaggle.com/c/GiveMeSomeCredit/ and German Credit (germancredit) for the additional experiments in Appendix C.
@@
It is based on the data from the 1994 US Census, where the main goal is to predict whether an individual earns more than 50,000\$ per year. The 14 features used to describe the instances include occupation, marital status, and education. Furthermore, it contains sensitive information such as age, gender, and race. In our experiments, we consider sensitive groups based on gender and race, as well as intersectional groups defined by their combination. This dataset is publicly available in the UCI repository. [^8] http://archive.ics.uci.edu/ml/index.php Although the dataset is originally divided into predefined training and test sets, we merge them and generate new train/test splits for our empirical evaluation. We pre-process the dataset as detailed in pawelczyk1carla.
@@
It includes financial and demographic information for 150,000 individuals and is used to predict the likelihood of experiencing financial distress within two years. Each record consists of 10 feature variables, such as age, monthly income, and credit indicators, and a binary target variable indicating whether the individual experienced a serious delinquency (90+ days past due) within that timeframe (0 = Yes, 1 = No). For our fairness-aware analysis, we used age as the sensitive attribute, dividing individuals into two groups: those aged 30 or younger, and those older than 30. This distinction reflects real socioeconomic and financial differences relevant to credit risk modeling. Younger individuals often lack established credit histories and income stability, whereas older individuals tend to have more financial experience and responsibilities, which can affect their credit behavior in meaningful ways. Moreover, in our experiments, we applied the following modifications:
@@
    We pre-processed the dataset following the procedure of detoni2024personalized to ensure recourse tractability.
@@
It collects information about several individuals from a German bank’s data from 1994. It contains details about the socioeconomic situation of individuals, such as their employment, housing, savings, etc. Besides, the set of features includes some sensitive information as well, such as gender and age. In this classification task, the objective is to predict whether an individual should obtain a good or bad credit score. This dataset is considerably smaller than the previous ones, containing only 1,000 instances and 20 features, and it is publicly available in the UCI repository. In our experiments, age has been considered as the sensitive attribute, and we have divided the instances into two (sensitive) groups based on whether their age is greater than 30 or less than or equal to 30. We pre-processed the dataset as detailed in pawelczyk1carla.